OpenStudy (anonymous):

The grades on the last history exam had a mean of 75%. Assume the population of grades on history exams is known to be normally distributed with a standard deviation of 7. What percent of students earn a score between 70% and 82%?

OpenStudy (kirbykirby):

Let \(X\) be the grades of the history exam. Then, you are asked to find \(P(70 < X < 82)\). WHat you can do is standardize X to get: \[X=\frac{X-\mu}{\sigma} \] and obtain: \[ P(70 < X < 82) = P\left( \frac{70-75}{7}<\frac{X-75}{7}<\frac{82-75}{7}\right)\\ =P(-0.71<Z<1)\] and then use a standard normal table to find the probability

OpenStudy (anonymous):

0.6285 0.6101 0.6024 0.5892

OpenStudy (anonymous):

So we don't know what x is?

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

Well X is a random variable. You standardize it as above to find a z-score. Then you find the probability using standard normal tables.

OpenStudy (anonymous):

I'm absolutely awful at this. Can you help me go through it step by step. I'm having a very hard time grasping the concept.

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

So you have X being certain values for the history grades. It has a mean of 75, and standard deviation 7. So, the normal curve will look smtg like this: